# ReFRAME or Remain: Unsupervised Lexical Semantic Change Detection with Frame Semantics

### v1

@@ 1 Introduction
Lexical semantic change detection is a well-established field in Natural Language Processing (NLP), with several shared tasks for different languages, amongst which the most popular one is SemEval 2020 Task 1 Schlechtweg et al. (2020). Most of the modern models are based on either static word embeddings like word2vec Mikolov et al. (2013) or ELMo Peters et al. (2018) or contextualised ones such as BERT Devlin et al. (2019) or S-BERT Reimers and Gurevych (2019), the latter currently achieving the state-of-the-art (SOTA) results Cassotti et al. (2023). Despite architectural differences, these methods share a common assumption which is the distributional hypothesis Firth (1957) and can capture some aspects of lexical meaning (e.g., Iacobacci et al., 2016; Lenci et al., 2022; Pilehvar and Camacho-Collados, 2019; Yenicelik et al., 2020). Using them for modelling semantic change would therefore be the logical next step.
@@
However, co-occurence is only one perspective of lexical semantics. Some other perspectives/frameworks are Behavioural Profiles Gries and Divjak (2009), Frame Semantics Fillmore (1982), Conceptual Metaphor Lakoff and Johnson (1981) (see Geeraerts (2010) for a comprehensive overview of lexical semantics), and some of them have also been used in diachronic lexical semantics and semantic change research (e.g., Geeraerts, 1983; Jansegers and Gries, 2020; Law, 2019; Pettersson-Traba, 2016; Sovran, 2004; Sullivan, 2007; Sweetser, 1990; Vais, 2022).
Exploiting Frame Semantics to detecting lexical semantic change is the focus of this paper. Consider the English noun plane: in the 20th century, after the birth of the modern airplane in 1903, the aircraft meaning emerged and became more dominant while the dimensional meaning became less dominant. This is reflected in the shift in the frame distributions of plane, where aircraft-related frames such as commerce_sell, departing, self_motion increase in frequency and dimensional-related frames such as identicality, part_whole decrease in frequency, as shown in Figure 1. Our main hypothesis is that the degree of semantic change of a word is reflected in changes in the frames that word participates in, either as the trigger or as a frame element. Note that trying to set a new SOTA is NOT our goal, instead we want to bring more attention to more simple, transparent techniques that utilise explicit linguistic information.
@@
  • Because the method is interpretable by design, we can decompose change scores into contributions of individual frames and inspect supporting contexts, enabling targeted qualitative analyses and understand how a word has changed (as in the plane example above) in contrast to the opaqueness of dense embedding models [^3] Note that while we acknowledge the growing field of explainable AI and its diverse range of techniques (e.g., Integrated Gradients Sundararajan et al. (2017), LIME Ribeiro et al. (2016) or SHAP Lundberg and Lee (2017), see surveys in Arrieta et al. (2019); Madsen et al. (2022)), these methods are often post-hoc, add methodological and computational overhead, and may not yield explanations that are straightforward to inspect linguistically. This makes our approach particularly convenient for linguistic and humanities research, where direct, feature-level inspection is desired.
@@ 2 Related work
A long-standing and influential tradition in diachronic semantics follow the Distributional Hypothesis Firth (1957). The rise of corpus linguistics and concordancing makes this idea a practical methodology. Researchers began tracking a target word’s prominent collocates and recurring patterns across corpora and later across time slices, often using association measures for identifying characteristic collocations Church and Hanks (1990); Sinclair (1991). This approach naturally extends to diachrony, where time-sliced collocational profiles reveal how highly associated collocates emerge, fade, or are replaced across periods. The prominence of this line of work is reflected by the many case studies (e.g., Alba-Salas, 2007; Pettersson-Traba, 2021), methods Anonymous_Author (2026); Garcia and García Salido (2019); McEnery et al. (2019) and tools Anthony (2014); Brezina et al. (2020); Kilgarriff et al. (2014); Jurish (2015); Jurish and Nieländer (2020) developed around it. With advances in computational infrastructure and modelling, attention increasingly shifted to a second realisation of the distributional hypothesis, where co-occurrence information is encoded as numerical vectors (embeddings). Early approaches built count-based co-occurrence matrices Bullinaria and Levy (2007); Gulordava and Baroni (2011); Schütze (1998). Later work introduced predictive neural embeddings that compress distributional information into dense vectors, with better performance in capturing semantics Baroni et al. (2014); Bojanowski et al. (2017); Levy and Goldberg (2014); Mikolov et al. (2013); Pennington et al. (2014). More recently, contextualised models extend this paradigm by modelling individual instances of a word Devlin et al. (2019). For semantic change problems, standard workflows split data into time slices, learn word representations per slice, align the spaces, and quantify change by comparing the aligned representations (e.g., via cosine distance or clustering) Tahmasebi et al. (2021); Tahmasebi and Dubossarsky (2023). As LSC research moved toward standardised evaluation with shared tasks and datasets, compact vector representations with a single change score became more attractive. As a result, most work in the area, including the current SOTA, follow this approach (e.g., Cassotti et al., 2023; Geeraerts et al., 2023; Giulianelli et al., 2020; Kim et al., 2014; Kulkarni et al., 2015; Kutuzov et al., 2018; Montariol et al., 2021; Tahmasebi and Dubossarsky, 2023; Tahmasebi et al., 2021).
@@
Other theory-driven frameworks have also been employed in Diachronic Semantics (e.g., Geeraerts, 1983; Jansegers and Gries, 2020; Law, 2019; Pettersson-Traba, 2016; Sovran, 2004; Sullivan, 2007; Sweetser, 1990; Vais, 2022). However, these approaches often yield descriptive outputs, making it hard to compare systematically across languages and datasets. Recently, efforts have been made to benchmark these theory-driven approach with standard shared tasks (e.g., Kutuzov et al., 2021; Ryzhova et al., 2021 on grammatical profiles; Tang et al., 2023 on sense distributions). Their strong performances on shared-task benchmarks (sometimes even outperform many embedding-based methods) inspired us to test the effectiveness of using Frame Semantics information in capturing word meaning change.
@@
Some practical weaknesses of Frame Semantics are 1) the the substantial time and manual work needed for building corpora, 2) the lack of coverage for different languages as frame inventories are built from particular languages and cultural/institutional settings, although efforts have been made for cross-lingual applicability Boas et al. (2025), 3) the lack of coverage for historical and out-of-domain data as many senses, lexical items, constructions are missing from the inventory Hartmann et al. (2017); Palmer and Sporleder (2010). These may explain why FrameNet has mainly been used in specific linguistics case studies Blank (1999, 2003); Georgakopoulos (2018); Koch (1999a, 2008, b); Law (2019); Vais (2022) and not in unsupervised computational approaches to LSC detection. As far as we know, our work is the first one to employ Frame Semantics for this task.
@@ 3 Data and tasks
Following SemEval 2020 Task 1 on Unsupervised Lexical Semantic Change Detection Schlechtweg et al. (2020), we formulate the problem as either binary classification or graded ranking, corresponding to subtasks 1 and 2. In subtask 1, given a set of target words, the system predicts whether each word gained or lost at least one sense between two time periods. In subtask 2, the system ranks target words by their degree of semantic change. Annotating semantic change is challenging because it requires assessing meaning differences across many usage instances from each period. A widely used protocol is DURel Schlechtweg et al. (2018), where annotators compare pairs of sentences containing a target word and judge whether the instances are used in the same sense or in different senses. These pairwise judgments are then aggregated into a final change score, either binary or continuous. This procedure was used to create the SemEval 2020 Task 1 dataset, the most widely used one to compared different LSC detection systems. It covers four languages: English, German, Latin, and Swedish, with manual annotations for both subtasks and provides a diachronic corpus for each language split into two time periods. Using this benchmark enables direct comparison of our approach with SOTA systems under a shared evaluation setup.
@@
However, our approach requires high-quality frame-semantic parses, which are not currently available for all languages. While English benefits from extensive FrameNet resources and several mature automatic frame-semantic parsers, we were unable to identify comparably reliable off-the-shelf parsers for the other 3 languages. For English, we used the pretrained Frame Semantic Transformer model Chanin (2023), trained on FrameNet 1.7 and reported to achieve SOTA performance. It is released as a Python library, making it convenient for using and reproducing the results. For German, although there are several FrameNet-like resources, they were all relatively dated Boas (2002); Burchardt et al. (2009, 2006); Erk and Padó (2006). The most recent resource and parser we found is PFN-DE Bick (2022). However, it is only accessible through a hosted interface that only accepts text files up to 10,000 characters and processing larger volumes requires contacting the provider for a paid processing agreement 86. This makes it impractical to parse 140 million token historical German corpora Schlechtweg et al. (2020) at scale compared to a locally runnable, openly distributed pipeline and raises concern about transparency and reproducibility. Moreover, it has only been evaluated on modern Twitter data, making its reliability on historical German uncertain. For Swedish, Svenskt frasnät 2.0 (SweFN) Dataset Språkbanken Text (2024) provides a FrameNet-style resource. The latest attempt to build a pretrained Swedish frame parser with SweFN used the same architecture as the English ones mentioned above, but with substantially lower performance Dannélls and Virk (2025); Dannélls et al. (2024). In a small sanity check with 30 random, diverse and basic sentences, we also observed 13 questionable frame predictions, such as Cutting for Jag såg en film igår‘I saw a movie yesterday’ (see the results in table 4 in the Appendix). The low F1 scores, coupled with its failure on such simple examples raise concern about the model’s reliability. For Latin, we did not find any public frame-semantic resources or pretrained parsers.
The lack of available and quality frame-semantic parsers for German, Swedish and Latin could conflate semantic change with parsing bias, threatening internal validity. We therefore restrict the experiment to English where strong pretrained parsers exist. We compensate for the lack of multilingual quantitative experiments with a detailed and thorough qualitative analysis. A quick description of the data is presented in Table 1. Note that although we evaluate only English, frame semantics is in principle language-agnostic Fillmore (1982) so the approach should generalise once suitable frame resources and parsers exist.
@@ 4.1 FrameNet parsing
Because we only need sentences containing the target lemma, we first split both the lemmatised and raw data into lemma-specific subcorpora. Since the lemmatised and raw files contain the same number of sentences in the same order, we do this by scanning the lemmatised file to select sentences that include the target lemma and then retrieving the corresponding sentences from the raw (token) file by index. We then frame-parsed them using the Frame transformer model (Chanin, 2023). The parser is available in two variants, base and small, with similar reported performance. It has also been noted that the models can occasionally identify a frame but return empty outputs. We therefore use base by default and fall back to small when it fails. If both models fail, we skip the sentence. Across 49,416 sentences, we triggered the fallback for 2,413 lemmatised sentences (4.9%) and 1,052 raw sentences (2.1%). Amongst these, both models failed (and the sentence was skipped) for 805 lemmatised sentences (1.6%) and 336 raw sentences (0.6%).
One technical limitation of the parser is that it only returns the raw string and not the location of the frame elements. This makes it impossible to locate the target lemma in the parsed token data (one target lemma may appear in different surface forms). We therefore use lemmatised corpora rather than the raw corpora. To assess the impact of using lemmatised rather than raw text, we compared (within each target word and time period) the frame distributions derived from the lemmatised and raw corpora using Jensen–Shannon Divergence Menéndez et al. (1997). The resulting JSD values (in the table 5 in the Appendix) are uniformly small (all < 0.15, most < 0.10), indicating that using lemmatised corpora introduces only minor changes to the induced frame distributions.
@@ 4.2 Basic procedure
To obtain the frame distributions, we collected all the frames that the target lemma is either the frame trigger or is amongst the frame elements. At the frame instance level, we avoid duplication to prevent inflated counts. Each frame instance is counted once even if 1) the target lemma is both the trigger and a frame element in the same frame, or 2) the target lemma occurs in multiple frame elements of the same frame. At the sentence level, if the target lemma participates in multiple frames within a sentence, all such frames are included. Likewise, multiple (even similar) frame instances within the same sentence are counted as separate occurrences.
We then experiment with 2 types of frame distributions [^4] We intentionally excluded the frame – trigger only distribution. This is because some lemma only appear in frame elements and not evoke any frame, making using frame – triggers only yield sparse or even empty distributions for some targets, causing JSD to be unstable or even undefined.: frame–elements only (FE) and frame–triggers + frame–elements (FTFE). To quantify the distributional changes across periods, we use Jensen–Shannon divergence (JSD) Menéndez et al. (1997). JSD is a symmetric, smoothed variant of Kullback–Leibler Divergence Kullback and Leibler (1951), and with log base 2 it is bounded between 0 and 1, making scores easy to interpret and comparable across targets. In addition, JSD can be decomposed into per-item contributions, allowing fine-grained attribution of which items (in this case, frames) drive the observed change. Prior work on grammatical profiles Kutuzov et al. (2021); Ryzhova et al. (2021) quantified change with cosine distance between time-specific vectors, but cosine compares vector geometry (angles) rather than directly comparing probability distributions.

### v2

@@ 1 Introduction
Lexical semantic change detection is well-established in Natural Language Processing (NLP), with different shared tasks for different languages, amongst which the most popular one is SemEval 2020 Task 1 Schlechtweg et al. (2020). Most modern systems are based on either static word embeddings like word2vec Mikolov et al. (2013) or ELMo Peters et al. (2018), or contextualised ones such as BERT Devlin et al. (2019) or S-BERT Reimers and Gurevych (2019), the latter currently achieving the state-of-the-art (SOTA) results Cassotti et al. (2023). Despite architectural differences, these methods share a common assumption: the distributional hypothesis Firth (1957) , and can capture some aspects of lexical meaning (e.g., Iacobacci et al., 2016; Lenci et al., 2022; Pilehvar and Camacho-Collados, 2019; Yenicelik et al., 2020). Using them for modelling semantic change would therefore be the logical next step.
@@
However, co-occurence is only one perspective of lexical semantics. There are other perspectives/frameworks such as Behavioural Profiles Gries and Divjak (2009), Frame Semantics Fillmore (1982), Conceptual Metaphor Lakoff and Johnson (1981) (see Geeraerts (2010) for a comprehensive overview of lexical semantics), and some of them have also been used in diachronic lexical semantics and semantic change research (e.g., Geeraerts, 1983; Jansegers and Gries, 2020; Law, 2019; Pettersson-Traba, 2016; Sovran, 2004; Sullivan, 2007; Sweetser, 1990; Vais, 2022).
Exploiting Frame Semantics to detecting lexical semantic change is the focus of this paper. Consider the English noun plane: in the 20th century, after the birth of the modern airplane in 1903, the aircraft meaning emerged and became more dominant while the dimensional meaning became less dominant. This is reflected in the shift in the frame distributions of plane, where aircraft-related frames such as commerce_sell, departing, self_motion increase in frequency and dimensional-related frames such as identicality, part_whole decrease in frequency, as shown in Figure 1. Our main hypothesis is that the degree of semantic change of a word is reflected in changes in the frames that word participates in, either as the frame trigger or as a frame element. Note that trying to set a new SOTA is NOT our goal, instead we want to bring more attention to more simple, transparent techniques that utilise explicit linguistic information.
@@
  • Because the method is interpretable by design, we can decompose change scores into contributions of individual frames and inspect supporting contexts, enabling targeted qualitative analyses to understand how a word has changed (as in the plane example above) in contrast to the opaqueness of dense embedding models [^3] While we acknowledge the growing field of explainable AI and its diverse range of techniques (e.g., Integrated Gradients Sundararajan et al. (2017), LIME Ribeiro et al. (2016) or SHAP Lundberg and Lee (2017), see surveys in Arrieta et al. (2019); Madsen et al. (2022)), these methods are often post-hoc, add methodological and computational overhead, and may not yield explanations that are straightforward to inspect linguistically. This makes our approach particularly convenient for linguistic and humanities research, where direct, feature-level inspection is desired..
@@ 2 Related work
A long-standing and influential tradition in diachronic semantics follows the Distributional Hypothesis Firth (1957). The rise of corpus linguistics and concordancing makes this idea a practical methodology. Researchers began tracking a target word’s prominent collocates and recurring patterns across corpora and later across time slices, often using association measures for identifying characteristic collocations Church and Hanks (1990); Sinclair (1991). This approach naturally extends to diachrony, where time-sliced collocational profiles reveal how highly associated collocates emerge, fade, or are replaced across periods. The prominence of this line of work is reflected by the many case studies (e.g., Alba-Salas, 2007; Pettersson-Traba, 2021), methods Anonymous_Author (2026); Garcia and García Salido (2019); McEnery et al. (2019) and tools Anthony (2014); Brezina et al. (2020); Kilgarriff et al. (2014); Jurish (2015); Jurish and Nieländer (2020) developed around it. With advances in computational infrastructure and modelling, attention increasingly shifted to a second realisation of the distributional hypothesis, where co-occurrence information is encoded as numerical vectors (embeddings). Early approaches built count-based co-occurrence matrices Bullinaria and Levy (2007); Gulordava and Baroni (2011); Schütze (1998). Later work introduced predictive neural embeddings that compress distributional information into dense vectors, with better performance in capturing semantics Baroni et al. (2014); Bojanowski et al. (2017); Levy and Goldberg (2014); Mikolov et al. (2013); Pennington et al. (2014). More recently, contextualised models extend this paradigm by modelling individual instances of a word Devlin et al. (2019). For semantic change problems, standard workflows split data into time slices, learn word representations per slice, align the spaces, and quantify change by comparing the aligned representations (e.g., via cosine distance or clustering) Tahmasebi et al. (2021); Tahmasebi and Dubossarsky (2023). As LSC research moved toward standardised evaluation with shared tasks and datasets, compact vector representations with a single change score became more attractive. As a result, most work in the area, including the current SOTA, follow this approach (e.g., Cassotti et al., 2023; Geeraerts et al., 2023; Giulianelli et al., 2020; Kim et al., 2014; Kulkarni et al., 2015; Kutuzov et al., 2018; Montariol et al., 2021; Tahmasebi and Dubossarsky, 2023; Tahmasebi et al., 2021).
@@
Other theory-driven frameworks have also been employed in Diachronic Semantics (e.g., Geeraerts, 1983; Jansegers and Gries, 2020; Law, 2019; Pettersson-Traba, 2016; Sovran, 2004; Sullivan, 2007; Sweetser, 1990; Vais, 2022). However, these approaches often yield descriptive outputs, making it difficult to compare systematically across languages and datasets. Recently, efforts have been made to benchmark these theory-driven approaches with standard shared tasks (e.g., Kutuzov et al., 2021; Ryzhova et al., 2021 on grammatical profiles; Tang et al., 2023 on sense distributions). Their strong performances on shared-task benchmarks (sometimes even outperform many embedding-based methods) inspired us to test the effectiveness of using Frame Semantics information in capturing word meaning change.
@@
Some practical weaknesses of Frame Semantics are:
  • The substantial time and manual work needed for building corpora.
  • The lack of coverage for different languages as frame inventories are built from particular languages and cultural/institutional settings, although efforts have been made for cross-lingual applicability Boas et al. (2025).
  • The lack of coverage for historical and out-of-domain data as many senses, lexical items, constructions are missing from the inventory Hartmann et al. (2017); Palmer and Sporleder (2010).
These may explain why FrameNet has mainly been used in specific linguistics case studies Blank (1999, 2003); Georgakopoulos (2018); Koch (1999a, 2008, b); Law (2019); Vais (2022) and not in unsupervised computational approaches to LSC detection. As far as we know, our work is the first one to employ Frame Semantics for this task.
@@ 3 Data and tasks
Following SemEval 2020 Task 1 on Unsupervised Lexical Semantic Change Detection Schlechtweg et al. (2020), we formulate the problem as either binary classification or graded ranking, corresponding to subtasks 1 and 2. In subtask 1, given a set of target words, the system predicts whether each word gained or lost at least one sense between two time periods. In subtask 2, the system ranks target words by their degrees of semantic change. Annotating semantic change is challenging because it requires assessing meaning differences across many usage instances from each period. A widely used protocol is DURel Schlechtweg et al. (2018), where annotators compare pairs of sentences containing a target word and judge whether the instances are used in the same sense or in different senses. These pairwise judgments are then aggregated into a final change score, either binary or continuous. This procedure was used to create the SemEval 2020 Task 1 dataset, the most widely used one to compared different LSC detection systems. It covers four languages: English, German, Latin, and Swedish, with manual annotations for both subtasks, and provides a diachronic corpus split into two time periods for each language. Using this benchmark enables direct comparison of our approach with SOTA systems under a shared evaluation setup.
@@
However, our approach requires high-quality frame-semantic parses, which are not currently available for all languages. While English benefits from extensive FrameNet resources and several mature automatic frame-semantic parsers, we were unable to find comparably reliable off-the-shelf parsers for the other 3 languages. For English, we used the pretrained Frame Semantic Transformer model Chanin (2023), which is trained on FrameNet 1.7 and is reported to achieve SOTA performance. It is released as a Python library, making it convenient for using and reproducing the results. For German, although there are several FrameNet-like resources, they were all relatively dated Boas (2002); Burchardt et al. (2009, 2006); Erk and Padó (2006). The most recent resource and parser we have found is PFN-DE Bick (2022). However, it is only accessible through a hosted interface that only accepts text files up to 10,000 characters, and processing larger volumes requires contacting the provider for a paid processing agreement 86. This makes it impractical to parse 140 million token historical German corpora Schlechtweg et al. (2020) at scale compared to a locally runnable, openly distributed pipeline, and raises concern about transparency and reproducibility. Moreover, it has only been evaluated on modern Twitter data, making its reliability on historical German uncertain. For Swedish, Svenskt frasnät 2.0 (SweFN) Dataset Språkbanken Text (2024) provides a FrameNet-style resource. The latest attempt to build a pretrained Swedish frame parser with SweFN used the same architecture as the English ones mentioned above, but with substantially lower performance Dannélls and Virk (2025); Dannélls et al. (2024). In a small sanity check with 30 random, diverse and basic sentences, we also observed 13 questionable frame predictions, such as Cutting for Jag såg en film igår‘I saw a movie yesterday’ (see the results in table 4 in the Appendix). The low F1 scores, coupled with its failure in such simple examples, raise concern about the reliability of the model. For Latin, we did not find any public frame-semantic resources or pretrained parsers.
Because of the lack of available and quality frame-semantic parsers for German, Swedish and Latin, we restrict the experiment to English where strong pretrained parsers exist. We compensate for the lack of multilingual quantitative experiments with a detailed qualitative analysis. A quick description of the English data is presented in Table 1. Note that although we evaluate only English, frame semantics is in principle language-agnostic Fillmore (1982) so the approach should generalise once suitable frame resources and parsers exist.
@@ 4.1 FrameNet parsing
Because we only need sentences containing the target lemma, we first split both the lemmatised and raw data into lemma-specific subcorpora. Since the lemmatised and raw files contain the same number of sentences in the same order, we do this by scanning the lemmatised file to select sentences that include the target lemma, then retrieving the corresponding sentences from the raw (token) file by index. We then frame-parsed them using the Frame transformer model Chanin (2023). The parser is available in two variants, base and small, with similar reported performances. It has also been noted that the models can occasionally identify a frame but return empty outputs. We therefore use base by default and fall back to small when it fails. If both models fail, we skip the sentence. Across 49,416 sentences, we triggered the fallback for 2,413 lemmatised sentences (4.9%) and 1,052 raw sentences (2.1%). Amongst these, both models failed (and the sentences were skipped) for only 805 lemmatised sentences (1.6%) and 336 raw sentences (0.6%).
One technical limitation of the parser is that it returns the frame elements as raw strings without locations. This makes it impossible to locate the target lemma in the parsed token data (one target lemma may appear in different surface forms). We therefore use lemmatised corpora rather than raw corpora. To assess the impact of doing so, we compared (within each target word and time period) the frame distributions derived from the lemmatised and raw corpora using Jensen–Shannon Divergence Menéndez et al. (1997). The resulting JSD values (in the table 5 in the Appendix) are uniformly small (all < 0.15, most < 0.10), indicating that using lemmatised corpora introduces only minor changes to the induced frame distributions.
@@ 4.2 Basic procedure
To obtain the frame distributions, we collected all the frames in which the target lemma is either the frame trigger or is amongst the frame elements. At the frame instance level, we avoid duplication to prevent inflated counts. Each frame instance is counted once even if 1) the target lemma is both the trigger and a frame element in the same frame, or 2) the target lemma occurs in multiple frame elements of the same frame. At the sentence level, if the target lemma participates in multiple frames within a sentence, all such frames are included. Similarly, multiple (even similar) frame instances within the same sentence are counted as separate occurrences.
We then experiment with 2 types of frame distributions [^4] We intentionally excluded the frame – trigger only distribution. This is because some lemma only appear in frame elements and not evoke any frame, making using frame – triggers only yield sparse or even empty distributions for some targets, causing JSD to be unstable or even undefined.: frame–elements only (FE) and frame–triggers + frame–elements (FTFE). To quantify the distributional changes across periods, we use Jensen–Shannon divergence (JSD) Menéndez et al. (1997). JSD is a symmetric, smoothed variant of Kullback–Leibler Divergence Kullback and Leibler (1951), and with log base 2 it is bounded between 0 and 1, making scores easy to interpret and comparable across targets. In addition, JSD can be decomposed into per-item contributions, allowing fine-grained attribution of which items (in this case, frames) drive the observed change. Prior work on grammatical profiles Kutuzov et al. (2021); Ryzhova et al. (2021) quantified changes with cosine distance between time-specific vectors, but cosine compares vector geometry (angles) rather than directly comparing probability distributions.